OpenStudy (anonymous):

soybean meal is 16% protein; cornmeal is 8% protein. How many pounds of each should be mixed together in order to get 320 pind mixture that is 11% protein? how many pounds of cornmeal? how many pounds of soybean meal?

OpenStudy (anonymous):

tough

OpenStudy (anonymous):

i know. :(

OpenStudy (amistre64):

s + c = 320 ; s = 320-c s(.16) + c(.08) = 320(.11) (320-c)(.16) + c(.08) = 320(.11)

OpenStudy (amistre64):

51.2 -.16c +.08c = 35.2 -35.2 = c(.16 -.08) ----------------------------- 16 = c(.08) 16/.08 = c 1600/8 = c = 200

OpenStudy (amistre64):

200 cornmeal 120 soybean ------------ 320 total

OpenStudy (anonymous):

AHHHH i see. Thank you!

OpenStudy (amistre64):

yep :)